OpenStudy (anonymous):

ANSWERS GET MEDALS If Nick has $2,500 to invest and he wants to invest in a program that compounds twice a year so that he will have $3,000 in 3 years, what rate of interest must he find for his investment to be successful?

OpenStudy (anonymous):

number of periods = 3*3 = 6 3000 = 2500(1 + r)^6 (1+r)^6 = 1.2 1+ r = 1.2 ^ 1/6 = 1.03085 r = 0.03085 3.085 %

OpenStudy (anonymous):

correction: number of periods = 3*2 = 6